OpenStudy (anonymous):

Verifying trig identities: sec(x+y)=sec x sec y/1-tan x tan y

OpenStudy (anonymous):

You can write it as: \[\sec(x+y) = \frac{1}{\cos(x+y)}\] And after that use: \[\cos(x+y) = \cos(x)\cdot \cos(y) - \sin(x) \cdot \sin(y)\]

OpenStudy (anonymous):

Can you just try once??

OpenStudy (anonymous):

Then can I divide by sin x sin y to get it on the bottom?

OpenStudy (anonymous):

You should divide by cos(x)cos(y) as seeing from what our RHS wants,...

OpenStudy (anonymous):

Please show me what are you doing? Otherwise we will not be in sync..

OpenStudy (anonymous):

It's ok I figured it out(: thanks

OpenStudy (anonymous):

If in problem then feel free to tell me..